1) Worth While Entertainment has an expected return of 11.90%, a beta of 1.15, and the expected return on the market is 10.90%. What must the risk-free rate be? Show calculations.

2) Sixth Fourth Bank has an issue of perpetual preferred stock with a $5.50 stated dividend that has a market price of $97 per share. What is the bank’s cost of preferred stock if the tax rate is 30%? Show calculations.
Answer #1

1)

Expected Return on Stock = Risk Free Rate+Beta[Return on Market-Risk Free Rate]

Therefore,

11.9 = R+1.15[10.9-R]

11.9 = R+12.535-1.15R

0.15R = 12.535-11.9 = 0.635

Therefore, Risk Free Rate = 0.635/0.15 = 4.23%

2)

Cost of Preferred Stock = Dividend/Market Price = 5.5/97 = 0.0567 = 5.67%

Note: Tax Rate is Irrelevant because Dividend is Not a Deductible expense.